# Multimodal Wireless Foundation Models

### v1

@@ I Introduction
Recent efforts demonstrated WFMs promising capabilities [rfm_gc, aboulfotouh2024building6gradiofoundation, wavesfm, wirelessgpt, lwm, contrastivefm, csi2vec, mashaal]. Early work in this area shows that self-supervised masked modeling can learn general wireless representations which transfer to downstream tasks [rfm_gc, aboulfotouh2024building6gradiofoundation]. WavesFM is pretrained with masked modeling to learn general and transferable representations for image-like wireless modalities (e.g., spectrograms, channel-state information (CSI), resource grids). It jointly performs wireless sensing, positioning and communication-related tasks, showing strong out-of-distribution generalization [wavesfm]. Wireless channel foundation models that operate on CSI and learn general-purpose embeddings for downstream tasks have also been explored in [wirelessgpt, lwm, contrastivefm, csi2vec]. Masked modeling has been used for CSI-based pretraining with downstream evaluation on several tasks: channel estimation, LOS/NLOS identification, and beam prediction [wirelessgpt, lwm]. Contrastive learning for CSI modeling has also been applied in [contrastivefm, csi2vec]. CSI2Vec adopts a triplet-based contrastive loss for pretraining and demonstrates downstream transfer to positioning and channel charting tasks [csi2vec]. For raw in-phase and quadrature (IQ) streams, IQFM is pretrained using a contrastive approach to learn general features for IQ-centric downstream tasks. It performs modulation classification, RF device fingerprinting, and angle-of-arrival estimation. It also shows high sample efficiency, adapting to new tasks with few-shot learning [mashaal].
A common limitation of current WFMs is they handle only one modality. Wireless systems, however, provide various observations (views) of the same over-the-air transmission, obtained at different stages of the physical layer stack, such as raw IQ samples, CSI, bits, and link quality measurements. Depending on the task and operating conditions, the most informative modality changes - and no single modality is best for all tasks. WFMs should therefore be designed to accept multiple modalities to enable a broader and more diverse range of tasks and scenarios. Potential benefits also include improved robustness via redundancy and complementary cues, as well as opportunities for cross-modal self-supervision, which has been shown to improve generalization [crossmodal, multimodalsurvey]. This motivates multimodal wireless foundation models (WFMs) that can natively handle multiple input modalities.
@@
The remainder of the paper is organized as follows. Section II first describes the wireless multimodal learning problem. Section III then presents the proposed ViT-based multimodal architecture, self-supervised pretraining data and methodology, the fine-tuning techniques along with the downstream tasks. Section IV presents the experimental results and analyzes performance across tasks and baselines. Section V summarizes the main findings and outlines directions for future work.
@@ 2 Initialize the loss value $\mathcal{L}\leftarrow 0$
where $\mathbf{E}_{\text{pos}}^{\mathrm{G}}\in\mathbb{R}^{N\times D_{\text{enc}}}$ and $\mathbf{E}_{\text{pos}}^{\mathrm{Q}}\in\mathbb{R}^{K\times D_{\text{enc}}}$ are 2D and 1D sinusoidal positional embeddings, respectively, and $(\mathbf{E}_{\text{ant}})_{k}=\mathbf{e}_{a_{k}}^{\top}.$ For details on sinusoidal positional embeddings, see [vit, wavesfm].
@@
Autoencoder Design. The encoder operates on the set of visible (unmasked) tokens, processing them through a series of ViT blocks to output feature tokens. These tokens are then projected to the decoder dimension, after which mask tokens are inserted at the previously masked positions. The mask token is a single learnable embedding shared across all masked positions, indicating where reconstruction is required. Positional embeddings are then added to the sequence to preserve ordering, which is then processed by the decoder through a series of ViT blocks. The design is asymmetric: most of the computation budget is in the encoder. This allows using the decoder for reconstruction during pretraining only and it can be safely discarded afterward. We do not show the details of ViT forward pass. For more details, refer to [vit, wavesfm].
@@ III-B Pretraining Data
Spectrogram Dataset. The dataset represents image-like wireless modalities. It is a collection of spectrograms for various signal types and bands (e.g., WiFi, LTE, Bluetooth, 5G-NR, and ISM). Recordings are captured over the air using SDRs at various center frequencies in the sub-6 GHz band, and sampling frequency in the range $\big[10,\ 60\big]$ MHz and converted to spectrogram representations; see [rfm_gc] for details of spectrogram computation. The dataset contains a total of 3200 samples.
IQ Dataset. The datasets represents the IQ modality. It consists of baseband IQ recordings from a 4-antenna MIMO setup in an indoor testbed, where signals with different modulations and technologies are exchanged between a transmitter and a receiver under various antenna configurations and transmitter/receiver locations; see [mashaal] for more details about the setup. The dataset contains a total of 3200 samples.
@@ III-C Fine-Tuning Configurations
Low-Rank Adaptation (LoRA) adapters. Low-rank adapters are inserted into the self-attention projections (query and value) of each encoder block. All original encoder weights remain frozen; trainable parameters are the LoRA adapters, the task head, and the input modality projections. This enables efficient adaptation while keeping the encoder shared among tasks.
@@ III-D Downstream Tasks
RF Fingerprinting (RFP). Raw IQ traces of LTE and 5G-NR transmissions collected at a fixed receiver from four different base stations. The target is device identification, i.e., which transmitter generated each segment [rfp].
Interference Detection (INTD) and Classification (INTC). Over-the-air LTE IQ recordings under clean conditions and under direct-sequence spread-spectrum (DSSS) interference. The task is to detect the presence of interference and, when present, identify its modulation type directly from IQ [icarus].
Human Activity Sensing (HAS). Wi-Fi CSI measurements collected while human subjects perform six activities between two access points with three antennas each. Input to the model is the CSI observation and the target is activity label [csi_sensing].
RF Signal Classification (RFS). Spectrogram measurements for various signal types, such as WiFi, FM radio, cellular, Bluetooth, and others, with a total of 20 classes. The task is to predict the signal type from its spectrogram [Zahid2024].
5G NR Positioning (POS). CSI measurements collected using sounding reference signals in a 5G NR setup. The target is to predict UE location from the CSI observations [5g_positioning].
@@ IV-B Baseline Downstream Performance
On IQ-centric tasks, the multimodal WFM consistently outperforms IQFM under both LP and FT2, with the exception of INTC under LP, showcasing strong generalization. Because IQFM is a compact model ($\sim 0.3$M parameters) compared to the multimodal WFM ($\sim 7$M), model capacity should be considered when reading the gaps. It is also reported in [mashaal] that IQFM significantly improves when fine-tuned with LoRA than with LP or partial finetuning.
@@ IV-C Multi-task Performance with LoRA
LoRA adapts the network by inserting low-rank modules, allowing both early and late blocks to change behavior while the backbone remains frozen. By placing these modules in attention and MLP layers throughout the encoder stack, both shallow and deep features can be adjusted. Compared to FT2, this is more flexible, keeps the encoder fully shared across tasks, and requires far fewer task-specific parameters. Our prior work showed that LoRA can exceed partial finetuning at a fraction of the parameter cost [wavesfm]. We fine-tune with LoRA rank $R=32$ and scaling $\alpha=32$, totaling $\sim\!0.3$M task-specific parameters, versus $\sim\!1.65$M for FT2. Results are reported in Table III.
@@
The multimodal WFM with LoRA is strong: it improves over FT2 on four of six tasks, with only small declines on HAS and INTD. Gain is most pronounced on RFS (+5%). LoRA achieves these results with $5\times$ fewer task-specific parameters, enabling full model sharing and, as a result, better multi-task capability.

### v2

@@ I Introduction
Recent efforts demonstrated WFMs promising capabilities [2, 1, 3, 15, 8, 10, 9, 5]. Early work in this area shows that self-supervised masked modeling can learn general wireless representations which transfer to downstream tasks [2, 1]. WavesFM is pretrained with masked modeling to learn general and transferable representations for image-like wireless modalities (e.g., spectrograms, channel-state information (CSI), resource grids). It jointly performs wireless sensing, positioning and communication-related tasks, showing strong out-of-distribution generalization [3]. Wireless channel foundation models that operate on CSI and learn general-purpose embeddings for downstream tasks have also been explored in [15, 8, 10]. Masked modeling has been used for CSI-based pretraining in [15]. Contrastive learning for CSI modeling has also been applied in [8, 10]. CSI2Vec adopts a triplet-based contrastive loss for pretraining and demonstrates downstream transfer to positioning and channel charting tasks [10]. For raw in-phase and quadrature (IQ) streams, IQFM is pretrained using a contrastive approach to learn general features for IQ-centric downstream tasks. It performs modulation classification, RF device fingerprinting, and angle-of-arrival estimation. It also shows high sample efficiency, adapting to new tasks with few-shot learning [9]. WirelessJEPA demonstrates pretraining on IQ streams using a joint-embedding predictive architecture, and shows generalization across six downstream tasks [5].
A common limitation of current WFMs is they handle only one modality. Wireless systems, however, provide various observations (views) of the same over-the-air transmission, obtained at different stages of the physical layer stack, such as raw IQ samples, CSI, bits, and link quality measurements. Depending on the task and operating conditions, the most informative modality changes - and no single modality is best for all tasks. WFMs should therefore be designed to accept multiple modalities to enable a broader and more diverse range of tasks and scenarios. Potential benefits also include improved robustness via redundancy and complementary cues, as well as opportunities for cross-modal self-supervision, which has been shown to improve generalization [11, 4]. This motivates multimodal wireless foundation models (WFMs) that can natively handle multiple input modalities.
@@
Reproducible research. Model weights, finetuning code and reproducibility instructions are available at: https://wavesfm.waveslab.ai.
@@ 2 Initialize the loss value $\mathcal{L}\leftarrow 0$
where $\mathbf{E}_{\text{pos}}^{\mathrm{G}}\in\mathbb{R}^{N\times D_{\text{enc}}}$ and $\mathbf{E}_{\text{pos}}^{\mathrm{Q}}\in\mathbb{R}^{K\times D_{\text{enc}}}$ are 2D and 1D sinusoidal positional embeddings, respectively, and $(\mathbf{E}_{\text{ant}})_{k}=\mathbf{e}_{a_{k}}^{\top}.$ For details on sinusoidal positional embeddings, see [6, 3].
@@
Autoencoder Design. The encoder operates on the set of visible (unmasked) tokens, processing them through a series of ViT blocks to output feature tokens. These tokens are then projected to the decoder dimension, after which mask tokens are inserted at the previously masked positions. The mask token is a single learnable embedding shared across all masked positions, indicating where reconstruction is required. Positional embeddings are then added to the sequence to preserve ordering, which is then processed by the decoder through a series of ViT blocks. The design is asymmetric: most of the computation budget is in the encoder. This allows using the decoder for reconstruction during pretraining only and it can be safely discarded afterward. We do not show the details of ViT forward pass. For more details, refer to [6, 3].
@@ III-B Pretraining Data
Spectrogram Dataset. The dataset represents image-like wireless modalities. It is a collection of spectrograms for various signal types and bands (e.g., WiFi, LTE, Bluetooth, 5G-NR, and ISM). Recordings are captured over the air using SDRs at various center frequencies in the sub-6 GHz band, and sampling frequency in the range $\big[10,\ 60\big]$ MHz and converted to spectrogram representations; see [2] for details of spectrogram computation. The dataset contains a total of 3200 samples.
IQ Dataset. The datasets represents the IQ modality. It consists of baseband IQ recordings from a 4-antenna MIMO setup in an indoor testbed, where signals with different modulations and technologies are exchanged between a transmitter and a receiver under various antenna configurations and transmitter/receiver locations; see [9] for more details about the setup. The dataset contains a total of 3200 samples.
@@ III-C Fine-Tuning Configurations
Low-Rank Adaptation (LoRA) adapters. Low-rank adapters are inserted into the query and value projections of each encoder block. All original encoder weights remain frozen; trainable parameters are the LoRA adapters, the task head, and the input modality projections. This enables efficient adaptation while keeping the encoder shared among tasks.
@@ III-D Downstream Tasks
RF Fingerprinting (RFP). Raw IQ traces of LTE and 5G-NR transmissions collected at a fixed receiver from four different base stations. The target is device identification, i.e., which transmitter generated each segment [12].
Interference Detection (INTD) and Classification (INTC). Over-the-air LTE IQ recordings under clean conditions and under direct-sequence spread-spectrum (DSSS) interference. The task is to detect the presence of interference and, when present, identify its modulation type directly from IQ [13].
Human Activity Sensing (HAS). Wi-Fi CSI measurements collected while human subjects perform six activities between two access points with three antennas each. Input to the model is the CSI observation and the target is activity label [14].
RF Signal Classification (RFS). Spectrogram measurements for various signal types, such as WiFi, FM radio, cellular, Bluetooth, and others, with a total of 20 classes. The task is to predict the signal type from its spectrogram [16].
5G NR Positioning (POS). CSI measurements collected using sounding reference signals in a 5G NR setup. The target is to predict UE location from the CSI observations [7].
@@ IV-B Baseline Downstream Performance
On IQ-centric tasks, the multimodal WFM consistently outperforms IQFM under both LP and FT2, with the exception of INTC under LP, showcasing strong generalization. Because IQFM is a compact model ($\sim 0.3$M parameters) compared to the multimodal WFM ($\sim 7$M), model capacity should be considered when reading the gaps. It is also reported in [9] that IQFM significantly improves when fine-tuned with LoRA than with LP or partial finetuning.
@@ IV-C Multi-task Performance with LoRA
LoRA adapts the network by inserting low-rank modules, allowing both early and late blocks to change behavior while the backbone remains frozen. By placing these modules in attention and MLP layers throughout the encoder stack, both shallow and deep features can be adjusted. Compared to FT2, this is more flexible, keeps the encoder fully shared across tasks, and requires far fewer task-specific parameters. Our prior work showed that LoRA can exceed partial finetuning at a fraction of the parameter cost [3]. We fine-tune with LoRA rank $R=32$ and scaling $\alpha=32$, totaling $\sim\!0.3$M task-specific parameters, versus $\sim\!1.65$M for FT2. Results are reported in Table III.
@@
The multimodal WFM with LoRA is strong: it improves over FT2 on four of five tasks, with only small declines on HAS and INTD. Gain is most pronounced on RFS (+5%). LoRA achieves these results with $5\times$ fewer task-specific parameters, enabling full model sharing and, as a result, better multi-task capability.